# Divertor shaping with neutral baffling as a solution to the tokamak power exhaust challenge

### v4

@@ 1 Introduction
This work shows the key experimental results of exploring novel divertor solutions on MAST-Upgrade. Since there is a continuum of ADC solutions, this work studies the impact of varying total flux expansion and divertor leg length instead of focusing only on the extreme Super-X topology, using plasmas with higher power (1.5-1.7 MW Neutral beam heating, $P_{SOL}=1.2$ MW). This not only provides the strongest experimental evidence to date for the benefits of ADCs by combining total flux expansion, divertor leg length and neutral baffling; but also shows these benefits are maintained at more moderate divertor shaping (lower $f_{R}$ and shorter leg lengths than the maximum values). These power exhaust benefits are obtained without any adverse core impact and include target heat flux reductions; improved access to, and stability of, plasma detachment; as well as improved core-edge compatibility. Since engineering complexity is a critical hurdle for integrating ADCs in reactors, the finding that their benefits can be maintained at more moderate shaping is of key importance for fusion reactor engineering and design [21, 20], advancing the path for using ADCs to achieve sustainable fusion energy.
@@ 2 Results
For each divertor configuration, the evolution of their power exhaust and detachment properties are diagnosed as the core electron density is gradually increased and the divertor conditions grow colder, whilst other parameters are held as constant as possible.
@@ 2.1 Benefits of combined strong neutral baffling, total flux expansion and divertor leg length
The longer-legged, totally flux expanded, divertors have a higher detachment stability to slow perturbations, qualitatively consistent with reduced models (section 4.3 [30]). The sensitivity of the detachment front to changes in core density, i.e, the slope of the detachment front position (Figure 2(b)) is a factor 5 steeper for the CD ($f_{GW}\approx 40\%$), compared to the ED and SXD at this core density: the detachment front is much more sensitive to changes in core density for the CD. These steady-state results indicate an inherent stabilisation, akin to a shock absorber, of the detachment front for the ED and SXD, in contrast to the CD where the ionisation region moves with minimal core changes out of the divertor chamber after detachment, increasing core and X-point radiation (Figure 2(h)). This suggests an enhanced detachment stability for the long-legged, totally flux expanded divertors. These benefits extend to dynamic variations in fuelling [38] and heating perturbations [36]: indicating an inherent lesser response of the detachment location to fuelling/heating transients and improved detachment control [38] for the ED and SXD.
The longer-legged, totally flux expanded, divertors result in larger heat flux reductions than expected. Based on magnetic geometry [27], a reduction in perpendicular heat flux by 5.8$\sim$ 5.8$\times$ 5.8 × and 2.1$\sim$ 2.1$\times$ 2.1 × for the SXD and ED is expected, compared to the CD, mainly due to increased poloidal ($F_{x}\sim 3.0\times$ (SXD) and 1.6$\sim$ 1.6$\times$ 1.6 × (ED)) and total ($F_{R}\sim 2.0\times$ (SXD) and 1.4$\sim$ 1.4$\times$ 1.4 × (ED)) flux expansion. However, a much larger reduction in target heat flux is observed: 18.5$\sim$ 18.5$\times$ 18.5 × and 7$\sim$ 7$\times$ 7 × for the SXD and ED, compared to the CD (Figure 2 c). The longer-legged divertors result in additional heat flux dissipation through volumetric and/or radial/cross-field transport by a factor 3.2$\sim$ 3.2$\times$ 3.2 ×, qualitatively consistent with both SOLPS-ITER simulations (section 2.3) [37] and volumetric power loss estimates (section 2.2.1).
@@ 2.2.1 Additional volume long-legged divertors drives power and particle losses
Particle balance analysis shows increased ion sinks reduce the target fluxes in the SXD and ED compared to the CD (3(a,b,c)), whereas the total ion source, inferred through particle balance, is the same within uncertainties for all three different geometries. Ion sinks are significant in both the SXD and the ED from the start of those discharges, both through MAR as well as EIR (in the SXD). Our spectroscopic analysis reveals plasma conditions of $n_{e}=2-4\times 10^{19}m^{-3}$ and $T_{e}\approx 0.2$ eV in the region where EIR becomes observable ($f_{GW}>33\%$ in the SXD and $f_{GW}>40\%$ in the ED) [36, 40]. MAR only appears in the CD at the highest core densities after its ionisation front detaches from the target ($f_{GW}>40\%$), but its magnitude remains limited downstream the baffle.
@@
Figure 4: The additional divertor leg length/volume of the SXD and ED configurations drives additional power and particle losses, whereas the ion sources/sinks and power losses are similar at the same poloidal distance to the X-point, according to the their 1D profiles along the divertor leg. Spectroscopically inferred line-integrated ion sources and sinks (part./$m^{2}$/s) for the SXD (a), ED (b) and CD (c) at $f_{GW}=35\%$ as function of poloidal distance to the X-point. The coloured dotted lines indicate their respective strike point positions, indicated by their magnetic geometry (d-f). (g) Power flow ($W$) towards the divertor targets as function of poloidal distance to the X-point from the divertor entrance to the target for the CD (red), ED (green) and SXD (blue) at $f_{GW}=35\%$, with a dotted line indicating their respective strike points. The part where the divertor leg is detached is indicated in grey. The power flow is inferred by subtracting from $P_{div}$ the cumulative sum of the hydrogenic power losses from upstream to the target.
The impact of divertor shaping on power and particle exhaust is revealed when studying the 1D profiles of ion sources and sinks (a-c), as well as power flows (g), along the divertor leg as function of poloidal distance to the X-point at a fixed core density ($f_{GW}=35\%$) (section 4). Both profiles are similar between the different geometries (up until the CD detachment onset) at the same poloidal distance to the X-point: the plasma is thus predominantly altered in the extended region. The deeper detachment and lower power loads in the SXD and CD are brought on by interactions in the additional volume available downstream of the ionisation region when the divertor leg is extended. Plasma-chemistry occurring in this region, resulting in MAR and MAD, plays a key role explaining the differences between the different divertor geometries.
2.2.2 Total flux expansion drives detachment onset reduction
@@
The impact of total flux expansion on the detachment onset predicted by the DLS model is consistent with our observations (section 2). The DLS predicted benefits for the SXD and ED over the CD arise mostly from an increase in total flux expansion. Given the density at which the CD detaches, the DLS predicts that the SXD and ED are already detached at the lowest core density achieved, consistent with the experiment. To compare the SXD and ED against each other, the onset of EIR is used as a colder reference point for DLS comparisons, showing agreement within 10 % of the experiment. Furthermore, the DLS model predicts that the detachment front position only depends on the magnetic field topology upstream of the detachment front (see Methods section), consistent with the observed invariance of the upstream parameters to the downstream divertor magnetic topology. That observation suggests that the additional leg length/shaping (e.g. total flux expansion) drives the observed differences between the SXD, ED and CD, consistent with DLS model predictions that only carries information on the divertor magnetic topology.
@@
Our power and particle balance analysis showed the additional volume in the SXD and ED is critical to explain the reduction of power and particle loads during detachment, whereas total flux expansion drives the improved access to detachment according to DLS model comparisons. The combination of increasing divertor leg length/volume and total flux expansion result in strong, synergistic, power exhaust benefits.
@@ 2.2.3 Neutral trapping enables exhaust benefits of total flux expansion and poloidal leg length
Although the poloidal divertor leg length downstream the baffle entrance is very different between the the different topologies, our experiments suggest that the baffling has a similarly strong impact on the CD (up until its detachment onset), ED and SXD. The observed and simulated (subsection 2.3) neutral pressures are similar between the three different divertor pressures. Likewise, the neutral trapping indicated by the ratio between ion source downstream the baffle ($|z|>1.3$ m Figure 1) to the total ion source [35], is similar between the three divertor topologies in simulations (56 % for the SXD and ED and 52 % for the CD, respectively). However, the increased detachment front sensitivity of the CD likely results in a reduction of neutral trapping after its detachment onset, diminishing the impact of neutral baffling on the CD after its detachment onset.
Strong neutral trapping, obtained by strong baffling, enables the shaping benefits of total flux expansion and poloidal leg length. Strong baffling augments plasma-neutral interactions in the detached volume by increasing the neutral density. This would increase energy and momentum transfer from elastic collisions to the plasma to the neutral cloud and results in the strong levels of MAR and MAD observed to drive power and particle losses in the SXD and ED (section 2.2.1), which are higher than those observed in long-legged open divertor geometries [43, 39, 8].
@@
The benefits of long-legged divertors, total flux expansion, divertor shaping and neutral baffling have been individually studied on TCV [29, 27]; showing benefits of neutral baffling [32, 44, 45, 46] and long divertor leg lengths [29]. The benefits of total flux expansion (both in terms of detachment onset and front sensitivity/stability) were, however, much smaller than predicted by the DLS model [47, 48]. SOLEDGE2D-EIRENE simulations suggest that the neutral baffling on TCV may be insufficient to recover the full benefit of total flux expansion [49], consistent with previous SOLPS-ITER simulations [35]. This can lead to strong plasma flows from the midplane to the target due to neutrals escaping the divertor [50, 51], which can diminish the impact of total flux expansion on the detachment onset [47]. Furthermore, SOLPS-ITER simulations [52] showed the openness of the TCV tokamak (without baffles), combined with the incidence angle of the divertor leg to the surface, resulted in a worse neutral trapping for the Super-X divertor than the conventional divertor. This not only negates part of the benefits of total flux expansion due to an ion flow from upstream towards the target, but also reduces the benefit of the neutrals in power/particle dissipation between the Super-X divertor compared to the conventional divertor.
MAST-U [53] uniquely integrates strong baffling and extreme divertor shaping, enabling strongly baffled novel divertor solutions. The DLS detachment onset predictions ($C_{t}$) varies by 110 % for the MAST-U CD to SXD strike point scan discharge. Comparatively, $C_{t}$ varies by 70 % for a range of TCV divertor geometries (target radius scan, poloidal flux expansion scan, as well as for the X-point target divertor) [27, 29]. By integrating strong baffling, long-legged divertors and total flux expansion, MAST-U enables retrieving the full benefit of its extreme shaping capability on the detachment onset and stability (factor 5×5$\times$5 × reduced sensitivity for ED & SXD vs CD, Figure 2(b)), as well as power and particle exhaust capability.
@@ 2.3 Strike point variations further confirm shaping benefits, consistent with exhaust simulations
The results in figure 2(b) suggested that the ionisation front position, once detached from the target, is invariant to the magnetic topology downstream of it: its location depends only on the upstream magnetic topology. This is confirmed by slowly sweeping the outer strike point from a CD to an ED to a SXD geometry at a constant core density (30 % Greenwald fraction, $n_{e}^{sep}\approx 0.8\times 10^{18}m^{-3}$) and power ($P_{SOL}\approx 1.0$ MW). The core conditions (not shown) remain mostly unchanged during this scan, with a 5-10% increase in upstream and core $T_{e}$ when transitioning from CD to SXD (potentially due to the longer connection length [4]). After the $D_{2}$ Fulcher band emission front, which is a proxy for the ionisation front [8, 54, 55], detaches from the target (target radius $\approx 0.95$ m), it remains close to this radial position as the strike point is swept further and further outwards and both total flux expansion and poloidal leg length is increased (Figure 5 (d-f)). This implies that the ionisation front position (for this $P_{SOL}$ and $f_{GW}$) remains at a constant poloidal distance to the X-point as the divertor leg length is further increased.
@@
Figure 5: Experimental data shows the ionisation region location is unaltered after detachment during a strike point scan, in quantitative agreement with SOLPS-ITER simulations. (a-c) Synthetic $D_{2}$ Fulcher emission from SOLPS-ITER simulations for the SXD (a), ED (b) and CD (c), overlaid with 5 eV contours (dotted lines) and the separatrix (solid line). (d-f) Experimentally measured $D_{2}$ Fulcher band emission (595-605 nm) for a strike point scan, moving from CD to SXD at constant density and power, obtained through inverting MWI imaging data for # 46895 [54]. (g-i) 1D ion sources and sinks, obtained from spectroscopic analysis integrated along the spectroscopic lines of sight (Figure 1b) (part./$m^{2}$/s), compared against synthetic diagnostic results from SOLPS-ITER simulations (dotted lines). For the SXD (g) two SOLPS-ITER simulation results are shown: one with default rates and one with corrected molecular charge exchange ($D_{2}+D^{+}\rightarrow D_{2}^{+}+D$) rates (’Sim. Corr. Rate’), obtained from [36], which increases MAR. To guide the eye, a magenta vertical line has been added at a radius of 0.95 m and an arrow has been added at the strike point location (a-i).
@@
This behaviour agrees with SOLPS-ITER predictions [37] of the CD, ED and SXD configurations (Figure 5 a-c). The CD simulation is attached, whereas the SXD and ED simulations are detached. The radius of both the $D_{2}$ Fulcher emission front as well as the 5 eV contour, for the ED and SXD, remains near $r=0.95$ m.
@@
A more detailed comparison between experiments and simulations is obtained by comparing their ion sources and sinks (Figure 5 (g-i)), indicating a quantitative agreement between experiments and simulations for the ion source and EIR. The MAR ion sinks are underestimated in the simulation in the detached region. This discrepancy is resolved when a corrected rate for molecular charge exchange is used in SOLPS-ITER, adopted from [36] (Figure 5 (g)).

### v5

@@ 1 Introduction
This work shows the key experimental results of exploring novel divertor solutions on MAST-Upgrade. Since there is a continuum of ADC solutions, this work studies the impact of varying total flux expansion and divertor leg length instead of focusing only on the extreme Super-X topology, using plasmas with higher power (1.5-1.7 MW Neutral Beam Injection (NBI) heating, $P_{SOL}=1.2$ MW). This not only provides the strongest experimental evidence to date for the benefits of ADCs by combining total flux expansion, divertor leg length and neutral baffling; but also shows these benefits are maintained at more moderate divertor shaping (lower $f_{R}$ and shorter leg lengths than the maximum values). These power exhaust benefits are obtained without any adverse core impact and include target heat flux reductions; improved access to, and stability of, plasma detachment; as well as improved core-edge compatibility. Since engineering complexity is a critical hurdle for integrating ADCs in reactors, the finding that their benefits can be maintained at more moderate shaping is of key importance for fusion reactor engineering and design [21, 20], advancing the path for using ADCs to achieve sustainable fusion energy.
@@ 2 Results
For each divertor configuration, the evolution of their power exhaust and detachment properties are diagnosed as the core electron density is gradually increased and the divertor conditions grow colder, whilst other parameters are held as constant as possible. For ease of reference and for comparison against literature, the line-averaged core electron density (obtained from interferometry) is expressed as a fraction to the maximum core density, Greenwald, limit [38] which depends on the plasma current and tokamak size.
@@ 2.1 Benefits of combined strong neutral baffling, total flux expansion and divertor leg length
The longer-legged, totally flux expanded, divertors have a higher detachment stability to quasi-steady-state changes in core parameters, qualitatively consistent with reduced (steady-state) models (section 4.3 [30]). The sensitivity of the detachment front to changes in core density, i.e, the slope of the detachment front position (Figure 2(b)) is a factor 5 steeper for the CD ($f_{GW}\approx 40\%$), compared to the ED and SXD at this core density: the detachment front is much more sensitive to changes in core density for the CD. These steady-state results indicate an inherent stabilisation, akin to a shock absorber, of the detachment front for the ED and SXD, in contrast to the CD where the ionisation region moves with minimal core changes out of the divertor chamber after detachment, increasing core and X-point radiation (Figure 2(h)). Although the reduced detachment sensitivity and increased operational window of detachment are related, they are not identical. Using the analogy of a shock absorber, a reduced detachment sensitivity corresponds to a stronger damping, whereas the wider detached operational regime increases the displacement the spring can undergo before the elastic limit is exceeded: both work in unison enhancing detachment stability for long-legged, totally flux expanded divertors.
These benefits extend to dynamic variations in fuelling [39] and heating perturbations [36]: indicating an inherent lesser response of the detachment location to fuelling/heating transients and improved detachment control [39] for the ED and SXD. In contrast to our quasi-steady-state experiments, [39] features sinusoidal core fuelling oscillations with more than 7 times faster fuelling changes to which the divertor responds dynamically.
The longer-legged, totally flux expanded, divertors result in larger heat flux reductions than expected. Based on the magnetic geometry [27], a reduction in perpendicular heat flux by 5.8$\sim$ 5.8$\times$ 5.8 × and 2.1$\sim$ 2.1$\times$ 2.1 × for the SXD and ED is expected, compared to the CD, mainly due to increased poloidal ($F_{x}\sim 3.0\times$ (SXD) and 1.6$\sim$ 1.6$\times$ 1.6 × (ED)) and total ($F_{R}\sim 2.0\times$ (SXD) and 1.4$\sim$ 1.4$\times$ 1.4 × (ED)) flux expansion. However, a much larger reduction in target heat flux is observed: 18.5$\sim$ 18.5$\times$ 18.5 × and 7$\sim$ 7$\times$ 7 × for the SXD and ED, compared to the CD (Figure 2 c). The longer-legged divertors result in additional heat flux dissipation through volumetric and/or radial/cross-field transport by a factor 3.2$\sim$ 3.2$\times$ 3.2 ×, qualitatively consistent with both SOLPS-ITER simulations (section 2.3) [37] and volumetric power loss estimates (section 2.2.1).
@@ 2.2.1 Additional volume long-legged divertors drives power and particle losses
Particle balance analysis shows increased ion sinks reduce the target fluxes in the SXD and ED compared to the CD (3(a,b,c)), whereas the total ion source is the same within uncertainties for all three different geometries. Ion sinks are significant in both the SXD and the ED from the start of those discharges, both through MAR as well as EIR (in the SXD). Our spectroscopic analysis reveals plasma conditions of $n_{e}=2-4\times 10^{19}m^{-3}$ and $T_{e}\approx 0.2$ eV in the region where EIR becomes observable ($f_{GW}>33\%$ in the SXD and $f_{GW}>40\%$ in the ED) [36, 41]. MAR only appears in the CD at the highest core densities after its ionisation front detaches from the target ($f_{GW}>40\%$), but its magnitude remains limited downstream the baffle.
The total ion source, inferred through particle balance, is obtained by adding the ion target flux and the ion sinks observed in the divertor chamber (see Methods section). This consists out of the divertor chamber ion source and any net inflow of ions into the divertor chamber. Previous work [36] showed the ion source is generated between the X-point and the ionisation front and up to 40% of it is generated in the divertor chamber. X-point imaging shows the neutral baffling is effective in limiting the ion source upstream of the X-point for all three configurations, consistent with simulation results (Figure 6, section 2.3).
@@
Figure 4: The additional divertor leg length/volume of the SXD and ED configurations drives additional power and particle losses, whereas the ion sources/sinks and power losses are similar at the same poloidal distance to the X-point, according to the their 1D profiles along the divertor leg. Spectroscopically inferred line-integrated ion sources and sinks (part./$m^{2}$/s) for the SXD (a), ED (b) and CD (c) at $f_{GW}=35\%$ as function of poloidal distance to the X-point. The coloured dotted lines indicate their respective strike point positions, indicated by their magnetic geometry (d-f). The 1D ion source/sink profiles (a,b,c) are extended downstream of their respective strike-points due to convolution of the radial-extent of the SOL/far-SOL with the spectroscopic lines-of-sight, where the plasma is colder than at the separatrix. (g) Power flow ($W$) towards the divertor targets as function of poloidal distance to the X-point from the divertor entrance to the target for the CD (red), ED (green) and SXD (blue) at $f_{GW}=35\%$, with a dotted line indicating their respective strike points. The part where the divertor leg is detached is indicated in grey. The power flow is inferred by subtracting from $P_{div}$ the cumulative sum of the hydrogenic power losses from upstream to the target.
The impact of divertor shaping on power and particle exhaust is revealed when studying the 1D profiles of ion sources and sinks (a-c), as well as power flows (g), along the divertor leg as function of poloidal distance to the X-point at a fixed core density ($f_{GW}=35\%$) (section 4). Both profiles are similar between the different geometries (up until the CD detachment onset) at the same poloidal distance to the X-point: the plasma is thus predominantly altered in the extended region. The deeper detachment and lower power loads in the SXD and ED are brought on by interactions in the additional volume available downstream of the ionisation region when the divertor leg is extended. Plasma-chemistry occurring in this region, resulting in MAR and MAD, plays a key role explaining the differences between the different divertor geometries.
2.2.2 Total flux expansion drives detachment onset and sensitivity reduction
@@
The impact of total flux expansion on the detachment onset predicted by the DLS model is consistent with our observations (table 2). The DLS predicted benefits for the SXD and ED over the CD arise mostly from an increase in total flux expansion. Given the density at which the CD detaches, the DLS predicts that the SXD and ED are already detached at the lowest core density achieved, consistent with the experiment. To compare the SXD and ED against each other, the onset of EIR is used as a colder reference point for DLS comparisons, showing agreement within 10 % of the experiment. Furthermore, the DLS model predicts that the detachment front position only depends on the magnetic field topology upstream of the detachment front (see Methods section), consistent with the observed invariance of the upstream parameters to the downstream divertor magnetic topology.
@@
Although only the magnetic geometry is used for obtaining DLS predictions on the differences between the SXD, ED and CD, it should be noted that the DLS model derivation assumes that: 1) the detachment front is infinitely thin; 2) all power dissipation is driven by impurity radiation. Both these assumptions are invalid for the MAST-U conditions shown. However, the agreement between the DLS model and the MAST-U results suggests that the impact of divertor topology on the detachment onset may be more generally applicable outside of impurity radiation dominant conditions. Further work is required generalising the DLS model for MAST-U like conditions, although the validity of these assumptions would increase in more reactor-relevant conditions.
Our power and particle balance analysis showed the additional volume in the SXD and ED is critical to explain the reduction of power and particle loads during detachment, whereas total flux expansion drives 80% of the improved access to detachment according to DLS model comparisons, with the longer divertor leg length driving the remaining 20%. The DLS model also predicts that total flux expansion reduces detachment sensitivity, qualitatively consistent with ED and SXD observations (benefit iii): the combination of increasing divertor leg length/volume and total flux expansion result in strong, synergistic, power exhaust benefits.
@@ 2.2.3 Neutral trapping enables exhaust benefits of total flux expansion and poloidal leg length
Although the poloidal divertor leg length downstream the baffle entrance is very different between the the different topologies, our experiments suggest that the baffling has a similarly strong impact on the CD (up until its detachment onset), ED and SXD. The observed and simulated (section 2.3) divertor neutral pressures are similar between the three geometries. Likewise, the neutral trapping, defined as the ratio between the ion source downstream the X-point to the total ion source [35], is similar between the three divertor topologies in simulations (78 % for the SXD and ED and 75 % for the CD, respectively), despite the SXD and ED configurations being deeply detached (section 2.3, Figure 6). In contrast, SOLPS-ITER simulations for a core density ramp of the (open, un-baffled) conventional TCV divertor indicated a neutral trapping of 32-45 % in detachment onset conditions, decreasing during deeper detachment to 11 % [45, 35]. However, the increased detachment front sensitivity of the CD likely results in a reduction of neutral trapping after its detachment onset, diminishing the impact of neutral baffling on the CD after its detachment onset.
@@
The benefits of long-legged divertors, total flux expansion, divertor shaping and neutral baffling have been individually studied on TCV [29, 27]; showing benefits of neutral baffling [32, 46, 47, 48] and long divertor leg lengths [29]. The benefits of total flux expansion (both in terms of detachment onset and front sensitivity/stability) were, however, much smaller than predicted by the DLS model [49, 50]. Escape of neutrals from the divertor to the SOL upstream of the X-point can lead to strong plasma flows from the midplane to the target [51, 52], which diminishes the impact of total flux expansion on the detachment onset [49]. SOLEDGE2D-EIRENE simulations suggest that the neutral baffling on TCV may be insufficient to recover the full benefit of total flux expansion [53], consistent with previous SOLPS-ITER simulations [35, 54] which showed that the neutral trapping of the (open, unbaffled) TCV Super-X divertor was worse than that of the conventional divertor. This not only negates part of the benefits of total flux expansion due to an ion flow from upstream the X-point towards the target, but also reduces the benefit of the neutrals in power/particle dissipation between the Super-X divertor compared to the conventional divertor.
This difference between MAST-U and TCV, as well as the absence of strong flows from the midplane to the X-point in MAST-U simulations, suggests that strong neutral trapping, obtained by baffling on MAST-U, enables the shaping benefits of total flux expansion. Additionally, neutral baffling augments power/momentum/particle losses from plasma-neutral interactions such as MAR and MAD by containing the neutrals in the divertor chamber, amplifying the benefits of long-legged divertors (section 2.2.1). Although neutral baffling is not required to obtain strong neutral trapping on high power devices (i.e., JET, AUG) near attached conditions, they would still be required to maintain high neutral trapping in cases where the ionisation source is significantly upstream of the target, motivating the STEP [55], SPARC [14] and ARC [17] divertor designs. The MAST-U results suggest that neutral baffling can be placed downstream the X-point: it is important that the baffle structure is: 1) upstream of the intended location of the ionisation front; 2) sufficiently prohibits the escape of neutrals to the SOL.
@@ 2.3 Strike point variations further confirm shaping benefits, consistent with exhaust simulations
MAST-U [56] uniquely integrates strong baffling and extreme divertor shaping, enabling strongly baffled novel divertor solutions. The DLS detachment onset predictions ($C_{t}$) varies by 110 % between MAST-U CD and SXD. Comparatively, $C_{t}$ varies by 70 % for a range of TCV divertor geometries (target radius scan, poloidal flux expansion scan, as well as for the X-point target divertor) [27, 29]. By integrating strong baffling, long-legged divertors and total flux expansion, MAST-U enables retrieving the full benefit of its extreme shaping capability on the detachment onset and stability (factor 5×5$\times$5 × reduced sensitivity for ED & SXD vs CD, Figure 2(b)), as well as power and particle exhaust capability. We will now investigate the results of a single discharge under constant constant core density (30 % Greenwald fraction, $n_{e}^{sep}\approx 0.8\times 10^{18}m^{-3}$) and power ($P_{SOL}\approx 1.0$ MW), where $C_{t}$ was altered by 110 % by slowly sweeping the outer strike point from a CD to an ED to a SXD geometry. No significant differences are observed, at the same strike point position and core density, during this strike point scan at constant core density with the density ramps presented previously.
@@
The results in figure 2(b) suggested that the ionisation front position, once detached from the target, is invariant to the magnetic topology downstream of it: its location depends only on the upstream magnetic topology. This is confirmed by the strike point sweep discharge, where the core conditions (not shown) remain mostly unchanged, with a 5-10% increase in upstream and core $T_{e}$ when transitioning from CD to SXD (potentially due to the longer connection length [4]). After the $D_{2}$ Fulcher band emission front, which is a proxy for the ionisation front [8, 57, 58], detaches from the target (target radius $\approx 0.95$ m), it remains close to this radial position as the strike point is swept further and further outwards and both total flux expansion and poloidal leg length is increased (Figure 5 (d-f)). This implies that the ionisation front position (for this $P_{SOL}$ and $f_{GW}$) remains at a constant poloidal distance to the X-point as the divertor leg length is further increased.
@@
Figure 5: Experimental data shows the ionisation region location is unaltered after detachment during a strike point scan, in quantitative agreement with SOLPS-ITER simulations. (a-c) Synthetic $D_{2}$ Fulcher emission from SOLPS-ITER simulations for the SXD (a), ED (b) and CD (c), overlaid with 5 eV contours (dotted lines) and the separatrix (solid line). (d-f) Experimentally measured $D_{2}$ Fulcher band emission (595-605 nm) for a strike point scan, moving from CD to SXD at constant density and power, obtained through inverting Multi-Wavelength-Imaging (MWI) imaging data for # 46895 [57]. (g-i) 1D ion sources and sinks, obtained from spectroscopic analysis integrated along the spectroscopic lines of sight (Figure 1b) (part./$m^{2}$/s), compared against synthetic diagnostic results from SOLPS-ITER simulations (dotted lines). For the SXD (g) two SOLPS-ITER simulation results are shown: one with default rates and one with corrected molecular charge exchange ($D_{2}+D^{+}\rightarrow D_{2}^{+}+D$) rates (’Sim. Corr. Rate’), obtained from [36], which increases MAR. To guide the eye, a magenta vertical line has been added at a radius of 0.95 m and an arrow has been added at the strike point location (a-i).
@@
This behaviour agrees with SOLPS-ITER predictions [37] of the CD, ED and SXD configurations (Figure 5 a-c). The CD simulation is attached, whereas the SXD and ED simulations are detached. The radius of both the $D_{2}$ Fulcher emission front as well as the 5 eV contour, for the ED and SXD, remains near $r=0.95$ m. The $D_{2}$ Fulcher emission near the X-point region is also in agreement between experiments (Figure 6), suggesting that the strong neutral trapping obtained in the simulations is consistent with the experiment.
However, the simulations feature an attached inner target (Figure 6 (a-c)) with strong $D_{2}$ Fulcher emission near the inner strike point (Figure 6 (d-f)), which is not observed experimentally (Figure 6 (g-i)). This suggests the inner target power loading is negligible in the experiments and overestimated in SOLPS-ITER simulations. This requires further study including using multi-diagnostic, Bayesian, analysis techniques [59] to infer plasma parameters outside the divertor chamber.
Figure 6: Overview of ion sources and $D_{2}$ Fulcher emission (as an ionisation location proxy [8, 57]) in the combined divertor and X-point region using simulations (a-f) for the SXD (a,d,g), ED (b,e,h) and CD (c,f,i) configurations, showing negligible ionisation upstream the X-point, consistent with experimental measurements (g-i). a-c) 2D ionisation source from SOLPS-ITER simulations (shown in Figure 5) with horizontal lines at $z=-1.6$ m (pink) and $z=1.07$ m (magenta), demarking the edge of the divertor spectroscopy view and X-point, respectively. The fraction of the ion source downstream these limits compared to the total ion source (outer leg only) are noted. d-f) Synthetic diagnostic for the $D_{2}$ Fulcher emissivity (arbitrary units) obtained from SOLPS-ITER simulations. g-i) Measured $D_{2}$ Fulcher emissivity (595-605 nm [57]) obtained from combined divertor imaging and X-point imaging inversions. The indicated time and discharges used are shown and are obtained from repeat discharges for the same core density as used in Figure 5 Only emissivities obtained at the same $r,z$ corresponding to the simulation grids are shown. An inversion artefact is present near $r=0.85$ m, $z=-1.6$ m, where there is a gap in coverage between the X-point and divertor imaging systems.
@@
A more detailed comparison between experiments and simulations is obtained by comparing their ion sources and sinks (Figure 5 (g-i)) in the outer divertor chamber, indicating a quantitative agreement between experiments and simulations for the ion source and EIR. The MAR ion sinks are underestimated in the simulation in the detached region. This discrepancy is resolved when a corrected rate for molecular charge exchange is used in SOLPS-ITER, adopted from [36] (Figure 5 (g)).